OpenStudy (anonymous):

Solve 4^(x + 7) = 6^(x – 1)

OpenStudy (jamesj):

You need to take the log of both sides, then you will have a linear equation in x which you can solve.

OpenStudy (jamesj):

For example on the left hand side ...\[\ln(4^{x+7}) = (x+7) \ln4\]

OpenStudy (jamesj):

Got it?

OpenStudy (anonymous):

so i need to switch the log to different sides?

OpenStudy (anonymous):

a different*

OpenStudy (anonymous):

side*

OpenStudy (jamesj):

No, just take the log of both sides and you see that you get terms like ln 4 and ln 6, which are just numbers.